# Topological frequency conversion in rhombohedral multilayer graphene

### v1

@@ Abstract
Figure 3: Valley-resolved frequency conversion ${\cal I}_{\rm FC}^{+}$ and dissipation ${\cal I}_{\rm diss}^{+}$ in bilayer graphene as a function of $\Omega=\sqrt{\omega_{1}\omega_{2}}$. The incidence angles $\theta_{j}$ are scaled to keep $\cos\theta_{j}/\omega_{j}$ constant, using Fig. 2 as a reference point. The dashed line $\sim\Omega^{2}$ corresponds to the low-frequency limit of ${\cal I}_{\rm FC}^{+}$, where non-adiabatic corrections can be neglected. We contrast (a) incommensurate and (b) commensurate frequency ratios $r=1/\sqrt{2}$ and $2/3$. For commensurate drives, frequency conversion overpowers dissipative losses in the THz regime.
@@
We first set $\omega_{1}\approx 0.8$ THz and probe different frequency ratios $r=\omega_{2}/\omega_{1}$. Fig. 2 shows the frequency conversion rates ${\cal P}_{\rm FC}^{\pm}(\mathbf{k})=\frac{1}{2}({\cal P}_{\rm 1}^{\pm}(\mathbf{k})-{\cal P}_{2}^{\pm}(\mathbf{k}))$, obtained for $r=1/\sqrt{2}$ (a–b) and $r={2}/{3}$ (d–e), respectively. For the latter case, we compute ${\cal P}_{\rm FC}^{\pm}(\mathbf{k})$ by averaging over $\Delta t$. Consistent with our analysis, electrons located within ${\cal A}_{\bm{\alpha}}$ transfer energy between the modes at the quantized rate $2hf_{1}f_{2}$. Near the boundary, these electrons generate substantial dissipation, ${\cal P}_{\rm diss}(\mathbf{k})=\sum_{\xi}({\cal P}_{\rm 1}^{\xi}(\mathbf{k})+{\cal P}_{2}^{\xi}(\mathbf{k}))$, through the Landau-Zener absorption mechanism [Fig. 2 (c,f)]. Dissipation is strongly reduced for commensurate driving due to the mechanism of Lissajous conversion [31]: Only a few electrons on the boundary of $\mathcal{A}_{\bm{\alpha}}$ are taken to the Weyl node by $e{\bf{A}}(t)/\hbar$ and hence generate Landau-Zener absorption. The anisotropic profile of conversion power along the boundary (panels d-e) arises from the distinct mode polarizations, which leads to a significant $\mathbf{k}$ dependence on the mode most responsible for pumping electrons between bands.
In Fig. 3 we plot the integrated conversion rate from a single valley, ${\cal I}_{\rm FC}^{+}\equiv\frac{1}{2}\left({\cal I}_{1}^{+}-{\cal I}_{2}^{+}\right)$, and the corresponding dissipated power ${\cal I}_{\rm diss}^{+}\equiv({\cal I}_{1}^{+}+{\cal I}_{2}^{+})$ as a function of frequency, with the same values of $r$ as in Fig. 2. To remove geometric factors from the analysis, we scale the incidence angle $\theta_{j}$ such that $\cos\theta_{j}/\omega_{j}$ remains constant, ensuring that the surface $\cal S_{\bm{\alpha}}$ is unchanged (electrons are only moving faster along it). In both cases, ${\cal I}_{\rm FC}^{+}$ initially grows quadratically with frequency, following the scaling of the conversion power quantum $hf_{1}f_{2}L$. This growth is eventually arrested due to the expansion of the boundary region of $\mathcal{A}_{\alpha}$, where electrons exhibit Landau-Zener absorption. This leads to a sub-quadratic scaling and, beyond a critical frequency, a sharp decrease in conversion power. Crucially, for commensurate driving on the THz scale [Fig. 3(b)], ${\cal I}_{\rm FC}^{\xi}$ can vastly outperform dissipative losses with $\tau$ of order $100-1000$ ps. In the Appendix we estimate typical heating rates of order $0.1-1$ K/ps, potentially allowing to perform topological frequency conversion over many THz drive cycles.
@@
Figure 4: Frequency conversion power (a,b) and dissipation (c,d) in valley $\xi=-1$ of rhombohedral trilayer graphene. We consider quasi-periodic (a,c) and periodic (b,d) driving, with parameters identical to Fig. 2. While electrons in the quantized regions contribute $50\%$ more to frequency conversion due to Weyl nodes carrying charge $L=3$, dissipation through Landau-Zener absorption is also enhanced because of the flatter low-energy dispersion.
@@
To demonstrate the extension of topological frequency conversion to other members of the RMG family, we simulate rhombohedral trilayer graphene irradiated by the same bichromatic driving as above (Fig. 4). Our results show a quantized plateau hosting topological frequency conversion at the rate $3hf_{1}f_{2}$ per electron (due to the charge-$3$ Weyl nodes in this system). This gain compared to bilayer graphene is however offset by a larger dissipation rate, due to a flatter low-energy dispersion that enhances the non-adiabatic region—a trend expected to continue with increasing number of layers.
@@
Our results indicate that amplification from a single valley can be achieved with commensurate frequencies in the THz range and relaxation timescales $\tau$ of order $100$ ps. In order to achieve net conversion, however, time-reversal symmetry (which relates the two valleys) must be broken. We speculate that valley-imbalanced populations can be realized either spontaneously through electronic interactions at low temperatures [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51], by interfacing with magnetic/spin-orbit coupled materials [52, 49, 53, 54, 55, 56, 57], or through photoexcitations generated by circularly-polarized light [58, 59, 60, 61, 62]. Recent experiments suggest that valley-imbalanced transient states can be long-lived (with valley lifetimes reported in a large range from $1\mu$s to $100$ms in recent bilayer graphene quantum dot experiments with applied magnetic fields [63, 64]). The investigation of such schemes present an interesting opportunity for further studies of frequency conversion phenomena.
@@
Finally, topological frequency conversion in RMG requires relatively strong electric fields and precise alignment of the polarization of the incoming radiation. We speculate that both challenges can be addressed by focusing the beams using waveguides and antennas—such structures, along with backgates of suitable geometry providing the required constant electric fields, could become part of a future frequency conversion device.
@@ Acknowledgements.
Appendix A Topological frequency conversion in slowly-driven two-level systems
@@
We now rewrite the time-dependent Hamiltonian Bloch vector ${\bf{d}}(t)$ as a function of the phases of the two modes, $\phi_{1}=\omega_{1}t$ and $\phi_{2}=\omega_{2}[t+\Delta t]$, such that ${\bf{d}}(t)=\mathbf{d}(\omega_{1}t,\omega_{2}t+\Delta\phi)$, where $\Delta\phi\equiv\omega_{2}\Delta t$. Then
$\partial_{t}\mathbf{d}(t)=[\omega_{1}\partial_{\phi_{1}}+\omega_{2}\partial_{\phi_{2}}]\mathbf{d}(\omega_{1}t,\omega_{2}t+\Delta\phi),$ (14)
with the phase difference $\Delta\phi=\omega_{2}\Delta t$. Inserting the above into Eq. (13), we recognize the contribution ${\bf{v}}\cdot\omega_{j}\partial_{\phi_{j}}\mathbf{d}$ as the rate of work done by mode $j$. The time-averaged conversion rate from mode $1$ to mode $2$ is given by $P_{\rm FC}=P_{1}=-P_{2}$ (neglecting dissipation), such that
@@
${\bf{v}}=\hat{\mathbf{d}}+\frac{\hbar}{2|\mathbf{d}|}\hat{\mathbf{d}}\times\partial_{t}\hat{\mathbf{d}}+\mathcal{O}\left(\frac{\omega^{2}}{|\mathbf{d}^{2}|}\right),$ (16)
where the second term on the r.h.s. denotes the leading-order diabatic correction and is ultimately responsible for topological frequency conversion. The result Eq. 16 can be derived from a sequence of rotating frame transformations, see e.g. Refs. [65, 36]. We can thus express ${\bf{v}}$ explicitly in terms of the phases $\phi_{1}$ and $\phi_{2}$,
${\bf{v}}=\hat{\mathbf{d}}+\frac{\hbar}{2|\hat{\mathbf{d}}|}\hat{\mathbf{d}}\times[\omega_{1}\partial_{\phi_{1}}+\omega_{2}\partial_{\phi_{2}}]\hat{\mathbf{d}}.$ (17)
@@
$\mathbf{v}\cdot\partial_{\phi_{1}}\mathbf{d}=\partial_{\phi_{1}}|\mathbf{d}|+\frac{\hbar}{2}(\hat{\mathbf{d}}\times[\omega_{1}\partial_{\phi_{1}}+\omega_{2}\partial_{\phi_{2}}]\hat{\mathbf{d}})\cdot\partial_{\phi_{1}}\hat{\mathbf{d}}$ (18)
where we also used that $(\hat{\mathbf{d}}\times\partial_{i}\hat{\mathbf{d}})\cdot\hat{\mathbf{d}}=0$. Next, we use the identity ${\bf{a}}\cdot({\bf{b}}\times{\bf{c}})={\bf{c}}\cdot({\bf{a}}\times{\bf{b}})$ along with the antisymmetry of the cross product to find
$\mathbf{v}\cdot\partial_{\phi_{1}}\mathbf{d}=\partial_{\phi_{1}}|\mathbf{d}|{-}\frac{\hbar\omega_{2}}{2}\hat{\mathbf{d}}\cdot(\partial_{{\phi_{1}}}\hat{\mathbf{d}}\times\partial_{{\phi_{2}}}\hat{\mathbf{d}}).$ (19)
We compute $P_{\rm FC}$ by averaging Eq. 19 over time. To this end, we restore explicit time (or phase) dependence and note that ${\bf{v}}(\phi_{1},\phi_{2})\cdot\partial_{\phi_{1}}\mathbf{d}(\phi_{1},\phi_{2})$ is $2\pi$-periodic in its arguments. For incommensurate frequencies, time-averaging is thus equivalent to phase-averaging:
@@
For commensurate frequencies, $P(\omega_{1}t,\omega_{2}t+\Delta\phi)$ is not identical to phase-averaging becuase the system only explores a one-dimensional, closed contour on the “phase Brillouin zone” $\Phi{\rm BZ}$ spanned by $\phi_{1},\phi_{2}$. However, we recover the incommensurate driving result when also averaging $P_{\rm FC}$ over the phase difference between the two modes $\Delta\phi$. (This procedure was used in our numerical simulations for commensurate driving presented in the main text.)
@@
$P_{\rm FC}=-\frac{\hbar\omega_{1}\omega_{2}}{2\pi}\int_{0}^{2\pi}\frac{{\rm d}\phi_{1}\phi_{2}}{4\pi}\hat{\mathbf{d}}\cdot(\partial_{\phi_{1}}\hat{\mathbf{d}}\times\partial_{\phi_{2}}\hat{\mathbf{d}}).$ (21)
@@
Appendix B Scaling of conversion power and dissipation with frequency
@@
Figure 5: Scaling of the (time-averaged) frequency conversion power ${\cal P}^{+}_{\rm FC}$ (panels a–e), and the corresponding dissipated power ${\cal P}^{+}_{\rm diss}$ (panels f–j), as a function of the geometric mean of incident frequencies, $\Omega=\sqrt{\omega_{1}\omega_{2}}$ (increasing from left to right). Color bars are expressed in units of the power quantum $hf_{1}f_{2}$. We consider incommensurate drives with $r=1/\sqrt{2}$.
@@
Figure 6: Scaling of the (time-averaged) frequency conversion power ${\cal P}^{+}_{\rm FC}$ (panels a–e), and the corresponding dissipated power ${\cal P}^{+}_{\rm diss}$ (panels f–j), as a function of the geometric mean of incident frequencies, $\Omega=\sqrt{\omega_{1}\omega_{2}}$ (increasing from left to right). Color bars are expressed in units of the power quantum $hf_{1}f_{2}$. We consider commensurate drives with $r=2/3$.
@@
In this Appendix we examine how the profile of frequency conversion and dissipation for electrons evolves as a function of frequency. As explained in the main text, we increase the driving frequencies $\omega_{j}$ in tandem with the incidence angles of the modes, $\theta_{j}$, in such a way that $\cos\theta_{j}/\omega_{j}$ remains constant. This ensures that the surface ${\cal S}_{\bm{\alpha}}$ along which the trajectories in $\bm{\alpha}$-space traversed by electrons remain fixed. In our analysis, we keep the ratio of frequencies $r=\omega_{1}/\omega_{2}$ fixed, and vary their geometric mean $\Omega=\sqrt{\omega_{1}\omega_{2}}$.
@@
We contrast incommensurate ($r=1/\sqrt{2}$, shown in Fig. 5) and commensurate ($r=2/3$, shown in Fig. 6) driving. In both cases, the quantized conversion plateau gradually shrinks with $\Omega$, due to the expansion of BZ areas dominated by non-adiabatic effects including, in particular, Landau-Zener absorption. The energy dissipation rate increases roughly linearly with frequency, and therefore becomes progressively less visible when normalized by the quantum of frequency conversion $hf_{1}f_{2}$.
@@
Periodic driving (Fig. 6) possesses two main advantages over quasi-periodic driving (Fig. 5), as discussed in the main text: frequency conversion persists at higher frequencies (panels a–e), and the total dissipated power is smaller (panels f–j). As a result, commensurate driving supports much stronger net conversion at THz-scale frequencies—as demonstrated in Fig. 3 of the main text.
@@
Appendix C Estimate of dissipation rate from nonzero chemical potential

### v2

@@ Abstract
Figure 3: Valley-resolved frequency conversion ${\cal I}_{\rm FC}^{+}$ and dissipation ${\cal I}_{\rm diss}^{+}$ intensities in Bernal bilayer graphene as a function of $\Omega=\sqrt{\omega_{1}\omega_{2}}$. The incidence angles $\theta_{j}$ are tuned simultaneously with $\Omega$ to keep the in-plane vector potential amplitude ($A_{j}\propto\cos\theta_{j}/\omega_{j}$) constant, using Fig. 2 as a reference point. The dashed lines indicate the scaling law ${\cal I}_{\rm FC}^{+}\sim\Omega^{2}$, which describes the low-frequency limit where non-adiabatic corrections can be neglected. The two panels contrast (a) incommensurate and (b) commensurate frequency ratios $r=1/\sqrt{2}$ and $2/3$, respectively. For commensurate drives, frequency conversion overpowers dissipative losses in the THz regime. Non-adiabatic effects cause the deviation of ${\cal I}_{\rm FC}^{+}$ from quadratic scaling at high frequencies; they are more prominent for incommensurate driving, leading to a vanishing conversion power for $\Omega\gtrsim 5\,{\rm THz}$ in (a).
@@
We first set $\omega_{1}\approx 0.8$ THz and probe different frequency ratios $r=\omega_{2}/\omega_{1}$. Fig. 2 shows the valley-resolved conversion power ${\cal P}_{\rm FC}^{\pm}(\mathbf{k})=\frac{1}{2}({\cal P}_{\rm 1}^{\pm}(\mathbf{k})-{\cal P}_{2}^{\pm}(\mathbf{k}))$, obtained for $r=1/\sqrt{2}$ (a–b) and $r={2}/{3}$ (d–e), respectively. For the latter case, we compute ${\cal P}_{\rm FC}^{\pm}(\mathbf{k})$ by averaging over $\Delta t$. Consistent with our analysis, electrons located within ${\cal A}_{\bm{\alpha}}$ transfer energy between the modes at the quantized rate ±$\pm 2hf_{1}f_{2}$. Near the boundary, these electrons generate substantial dissipation, ${\cal P}_{\rm diss}(\mathbf{k})=\sum_{\xi}({\cal P}_{\rm 1}^{\xi}(\mathbf{k})+{\cal P}_{2}^{\xi}(\mathbf{k}))$, through the Landau-Zener absorption mechanism [Fig. 2 (c,f)]. Dissipation is strongly reduced for commensurate driving due to the mechanism of Lissajous conversion [31]: Only a few electrons on the boundary of $\mathcal{A}_{\bm{\alpha}}$ are taken to the Weyl node by $e{\bf{A}}(t)/\hbar$ and hence generate Landau-Zener absorption. The anisotropic profile of conversion power along the boundary arises from the distinct mode polarizations, which lead to a significant $\mathbf{k}$ dependence on the mode most responsible for pumping electrons between bands.
@@
Figure 4: Frequency conversion power (a,b) and dissipation (c,d) in valley $\xi=+1$ of rhombohedral trilayer graphene. We consider quasi-periodic (a,c) and periodic (b,d) driving, with parameters identical to Fig. 2. While electrons in the quantized regions contribute $50\%$ more to frequency conversion than in bilayer graphene, due to Weyl nodes carrying charge $L=3$, dissipation through Landau-Zener absorption is also enhanced because of the flatter low-energy dispersion.
In Fig. 3 we plot the integrated conversion intensity from a single valley, ${\cal I}_{\rm FC}^{+}\equiv\frac{1}{2}\left({\cal I}_{1}^{+}-{\cal I}_{2}^{+}\right)$, and the corresponding dissipated intensity ${\cal I}_{\rm diss}^{+}\equiv({\cal I}_{1}^{+}+{\cal I}_{2}^{+})$ as a function of frequency, with the same values of $r$ as in Fig. 2. To remove geometric factors from the analysis, we scale the incidence angle $\theta_{j}$ such that the in-plane vector potential amplitude, $A^{\parallel}_{j}\propto E_{j}\cos\theta_{j}/\omega_{j}$, remains constant—thus the surface $\cal S_{\bm{\alpha}}$ is unchanged and only the time-dependence of $\bm{\alpha}(t)$ is rescaled. In both commensurate and incommensurate cases, ${\cal I}_{\rm FC}^{+}$ grows quadratically with $\Omega$ at low frequencies, following the scaling of the conversion power quantum $hf_{1}f_{2}L$. At higher frequencies, the expansion of the region with strong Landau-Zener tunneling, near the boundary of $\cal A_{\bm{\alpha}}$, leads to a sub-quadratic frequency scaling of ${\cal I}_{\rm FC}^{+}$ (see also Figs. S1 and S2 in the Appendix). These non-adiabatic effects are much stronger for incommensurate driving protocols [Fig. 3(a)]; beyond a critical frequency they lead to a sharp decrease in conversion power. Crucially, for commensurate driving on the THz scale [Fig. 3(b)], ${\cal I}_{\rm FC}^{\xi}$ can vastly outperform dissipative losses with $\tau$ of order $100-1000$ ps. In the Appendix we also estimate typical heating rates of order $0.1-1$ K/ps, potentially allowing to perform topological frequency conversion over many THz drive cycles.
@@
To demonstrate the extension of topological frequency conversion to other members of the RMG family, we simulate rhombohedral trilayer graphene irradiated by the same bichromatic driving as above (Fig. 4). Our results show quantized plateaus hosting topological frequency conversion at the rate ±$\pm 3hf_{1}f_{2}$ per electron (due to the charge-$3$ Weyl nodes in this system). This gain compared to bilayer graphene is however offset by a larger dissipation rate, due to the flatter low-energy dispersion that enhances the non-adiabatic regions of momentum space—a trend expected to continue with increasing number of layers [35, 36].
@@
Our results indicate that amplification from a single valley can be achieved with commensurate frequencies in the THz range and relaxation timescales $\tau$ of order $100-1000$ ps. In order to achieve a net conversion, however, time-reversal symmetry (which relates the two valleys) must be broken. We speculate that valley-imbalanced populations can be realized either spontaneously through electronic interactions at low temperatures [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55], by interfacing with magnetic/spin-orbit coupled materials [56, 51, 57, 52, 53, 54, 55, 58, 59, 60, 61, 62], or through photoexcitations generated by circularly-polarized light [63, 64, 65, 66, 67]. Recent experiments suggest that valley-imbalanced transient states can be long-lived (with, e.g., valley lifetimes reported in a large range from $1\mu$s to $100$ms in bilayer graphene quantum dot experiments with applied magnetic fields [68, 69]). The investigation of such schemes present an interesting opportunity for further studies of frequency conversion phenomena.
@@
Finally, topological frequency conversion in RMG requires relatively strong electric fields and precise alignment of the polarization of the incoming radiation. We speculate that these challenges can be addressed by focusing the beams using waveguides and antennas—such structures, along with backgates of suitable geometry providing the required constant electric fields, could become part of a future frequency conversion device.
@@ Acknowledgements.
Appendix A SUPPLEMENTARY MATERIAL
for
Topological frequency conversion in rhombohedral graphene multilayers
Étienne Lantagne-Hurtubise, Iliya Esin, Gil Refael, and Frederik Nathan
Appendix B Topological frequency conversion in slowly-driven two-level systems
@@
We now rewrite the time-dependent Hamiltonian Bloch vector ${\bf{d}}(t)$ as a function of the phases of the two modes, $\phi_{1}=\omega_{1}t$ and $\phi_{2}=\omega_{2}[t+\Delta t]$, such that ${\bf{d}}(t)=\mathbf{d}(\omega_{1}t,\omega_{2}t+\Delta\phi)$, with the phase difference $\Delta\phi\equiv\omega_{2}\Delta t$. Then
$\partial_{t}\mathbf{d}(t)=[\omega_{1}\partial_{\phi_{1}}+\omega_{2}\partial_{\phi_{2}}]\mathbf{d}(\omega_{1}t,\omega_{2}t+\Delta\phi).$ (13)
Inserting the above into Eq. (12), we recognize the contribution ${\bf{v}}\cdot\omega_{j}\partial_{\phi_{j}}\mathbf{d}$ as the rate of work done by mode $j$, $P_{j}$. The time-averaged conversion power from mode $1$ to mode $2$ is given by $P_{\rm FC}=P_{1}=-P_{2}$ (neglecting dissipation), such that
@@
${\bf{v}}=-\hat{\mathbf{d}}+\frac{\hbar}{2|\mathbf{d}|}\hat{\mathbf{d}}\times\partial_{t}\hat{\mathbf{d}}+\mathcal{O}\left(\frac{\omega^{2}}{|\mathbf{d}^{2}|}\right),$ (15)
where the second term on the r.h.s. denotes the leading-order diabatic correction and is ultimately responsible for topological frequency conversion. The result Eq. 15 can be derived from a sequence of rotating frame transformations, see e.g. Refs. [70, 38]. We neglect $\mathcal{O}(\omega^{2}/|{\bf{d}}|^{2})$ subleading nonadiabatic corrections in the following. In this case, we can express ${\bf{v}}$ explicitly in terms of the phases $\phi_{1}$ and $\phi_{2}$,
${\bf{v}}\approx-\hat{\mathbf{d}}+\frac{\hbar}{2|\mathbf{d}|}\hat{\mathbf{d}}\times[\omega_{1}\partial_{\phi_{1}}+\omega_{2}\partial_{\phi_{2}}]\hat{\mathbf{d}}.$ (16)
@@
$\mathbf{v}\cdot\partial_{\phi_{1}}\mathbf{d}\approx-\partial_{\phi_{1}}|\mathbf{d}|+\frac{\hbar}{2}\left(\hat{\mathbf{d}}\times[\omega_{1}\partial_{\phi_{1}}+\omega_{2}\partial_{\phi_{2}}]\hat{\mathbf{d}}\right)\cdot\partial_{\phi_{1}}\hat{\mathbf{d}}$ (17)
where we also used that $(\hat{\mathbf{d}}\times\partial_{i}\hat{\mathbf{d}})\cdot\hat{\mathbf{d}}=0$. Next, we use the identity $({\bf{a}}\times{\bf{b}})\cdot{\bf{c}}=({\bf{c}}\times{\bf{a}})\cdot{\bf{b}}$ along with the antisymmetry of the cross product to find
$\mathbf{v}\cdot\partial_{\phi_{1}}\mathbf{d}=-\partial_{\phi_{1}}|\mathbf{d}|-\frac{\hbar\omega_{2}}{2}\hat{\mathbf{d}}\cdot(\partial_{{\phi_{1}}}\hat{\mathbf{d}}\times\partial_{{\phi_{2}}}\hat{\mathbf{d}}).$ (18)
We compute $P_{\rm FC}$ by averaging Eq. (18) over time. To this end, we restore explicit time (or phase) dependence and note that ${\bf{v}}(\phi_{1},\phi_{2})\cdot\partial_{\phi_{1}}\mathbf{d}(\phi_{1},\phi_{2})$ is $2\pi$-periodic in its arguments. For incommensurate frequencies, time-averaging is thus equivalent to phase-averaging:
@@
For commensurate frequencies, $P(\omega_{1}t,\omega_{2}t+\Delta\phi)$ is not identical to phase-averaging because the system only explores a one-dimensional, closed contour on the “phase Brillouin zone” spanned by $\phi_{1},\phi_{2}$. However, we recover the incommensurate driving result when also averaging $P_{\rm FC}$ over the phase difference $\Delta\phi$ between the two modes. This procedure was used in our numerical simulations for commensurate driving presented in the main text.
@@
$P_{\rm FC}\approx-\frac{\hbar\omega_{1}\omega_{2}}{2\pi}\int_{0}^{2\pi}\frac{{\rm d}\phi_{1}\phi_{2}}{4\pi}\hat{\mathbf{d}}\cdot(\partial_{\phi_{1}}\hat{\mathbf{d}}\times\partial_{\phi_{2}}\hat{\mathbf{d}}).$ (20)
@@
Appendix C Scaling of conversion power and dissipation with frequency
@@
Fig. S1: Scaling of the (time-averaged) frequency conversion power ${\cal P}^{+}_{\rm FC}$ (panels a–e), and the corresponding dissipated power ${\cal P}^{+}_{\rm diss}$ (panels f–j), as a function of the geometric mean of incident frequencies, $\Omega=\sqrt{\omega_{1}\omega_{2}}$ (increasing from left to right). Color bars are expressed in units of the power quantum $hf_{1}f_{2}$. We consider incommensurate drives with $r=1/\sqrt{2}$.
@@
Fig. S2: Scaling of the (time-averaged) frequency conversion power ${\cal P}^{+}_{\rm FC}$ (panels a–e), and the corresponding dissipated power ${\cal P}^{+}_{\rm diss}$ (panels f–j), as a function of the geometric mean of incident frequencies, $\Omega=\sqrt{\omega_{1}\omega_{2}}$ (increasing from left to right). Color bars are expressed in units of the power quantum $hf_{1}f_{2}$. We consider commensurate drives with $r=2/3$.
@@
In this Appendix we examine how the profile of frequency conversion and dissipation for electrons evolves as a function of frequency. As explained in the main text, we increase the driving frequencies $\omega_{j}$ in tandem with the incidence angles of the modes, $\theta_{j}$, in such a way that $\cos\theta_{j}/\omega_{j}$ remains constant. This ensures that the surface ${\cal S}_{\bm{\alpha}}$ in $\bm{\alpha}$-space, traversed by electrons along their trajectories, remains fixed. In our analysis, we keep the ratio of frequencies $r=\omega_{1}/\omega_{2}$ constant, and vary their geometric mean $\Omega=\sqrt{\omega_{1}\omega_{2}}$.
@@
We contrast incommensurate ($r=1/\sqrt{2}$, Fig. S1) and commensurate ($r=2/3$, Fig. S2) driving. In both cases, the quantized conversion plateaus gradually shrink with $\Omega$, due to the expansion of Brillouin zone areas dominated by non-adiabatic effects including, in particular, Landau-Zener absorption. We find that the energy dissipation rate increases roughly linearly with frequency, and therefore becomes progressively less visible when normalized by the quantum of frequency conversion $hf_{1}f_{2}$.
@@
Periodic driving (Fig. S2) possesses two main advantages over quasi-periodic driving (Fig. S1), as discussed in the main text: frequency conversion persists at higher frequencies (panels a–e), and the total dissipated power is smaller (panels f–j). As a result, commensurate driving supports much stronger net conversion at THz-scale frequencies—as demonstrated in Fig. 3 of the main text.
@@
Appendix D Estimate of dissipation rate with a non-zero chemical potential